OpenStudy (sonyalee77):

For the equation -2(x+1) = 8, Caitlin distributes the 2 to get, -2x +2 = 8. Then she subtracts the 2 from both sides to get -2x = 6. Lastly she divides both sides by -2 to get x =-3. Did Caitlin use the distributive property correctly?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

its not plus 2, its minus 2

OpenStudy (anonymous):

-2x-2=8

OpenStudy (anonymous):

-2x-2=8 -2x=10 etc

OpenStudy (anonymous):

\[-2(x+1)=8\] becomes \[-2x-2=8\] becomes \[-2x=10\] becomes \[x=10/-2\]

OpenStudy (anonymous):

so no its -5

OpenStudy (sonyalee77):

So she didn't use the distributive property right"

OpenStudy (anonymous):

Caitlin did i wrong

OpenStudy (anonymous):

did it

OpenStudy (sonyalee77):

Okay thanx a whole lot!

OpenStudy (anonymous):

not solving for x, she still did not use the distributive property correctly. like the first few replies you're suppose to multiply the number with each number in the parenthesis so -2(x+1) essentially equals to (-2 * x) + (-2 * 1) which gets you -2x-2 .